OpenStudy (anonymous):

raise the quantity in parenthesis to the indicated exponent, and simplify the resulting expression. Express with positive exponents. (43x^-4y^4/86x^-3y^-5)^3

OpenStudy (anonymous):

\[(43x ^{-4}y ^4 / 86x ^{-3}y ^{-5})^{3}\]

OpenStudy (anonymous):

so i have gotten this far.....\[1x ^{-12}y ^{12}/2x ^{3}y ^{-15}\]

OpenStudy (shane_b):

It's a bit tricky...let's hope I don't make a silly mistake typing it all out :)\[(\large \frac{43x ^{-4}y ^4}{86x ^{-3}y ^{-5}})^{3}\]\[\large (\frac{y^4y^5x^3}{2x^4})^3\]\[\large (\frac{y^9}{2x^4x})^3\]\[\large \frac{y^{27}}{8x^{15}}\]

OpenStudy (anonymous):

ok, ty what happens to the 43/86 ?

OpenStudy (shane_b):

It turned into 1/2....I just dropped the 1 from the top

OpenStudy (anonymous):

oh ok, gotcha! thanks for your help! algebra is so not my forte!! lol

OpenStudy (shane_b):

It's ok...let me check it on paper...just to be sure

OpenStudy (anonymous):

thanks again, I am in my 2nd algebra class for college and I barely passed the first one :)

OpenStudy (shane_b):

There is a mistake :)

OpenStudy (anonymous):

okie doke. I'm writing it out too

OpenStudy (shane_b):

\[(\large \frac{43x ^{-4}y ^4}{86x ^{-3}y ^{-5}})^{3}\]\[\large (\frac{y^4y^5x^3}{2x^4})^3\]\[\large (\frac{y^9}{2x})^3\]\[\large \frac{y^{27}}{8x^{3}}\]

OpenStudy (shane_b):

I made a silly mistake with the x's....but this should be correct now.

OpenStudy (anonymous):

thanks a bunch!

OpenStudy (shane_b):

np :)

OpenStudy (shane_b):

BTW: If something is unclear, just ask

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

ok,so inthe second to last step, where did the 4 go in 2x^4 ?

OpenStudy (shane_b):

I just divided and simplified it along the way:\[\large \frac{x^3}{x^4}=x^{-1}=\frac{1}{x}\]

OpenStudy (anonymous):

gotcha ok ty

OpenStudy (shane_b):

You're welcome. This all just takes a little practice then it gets easier.

OpenStudy (anonymous):

I hope so lol my little brother is way better at this stuff and he's 7yrs younger than me :)